OpenStudy (anonymous):

x2 + 4x – 12 = 0 by completing the square.

OpenStudy (anonymous):

add 12 to both sides \[x^2+4x=12\] take half of 4 which is 2 and write \[(x+2)^2=12+2^2=12+4=16\] \[(x+2)^2=16\] take the square root and get \[x+2=4\] or \[x+2=-4\] so \[x=2\] or \[x=-6\]

OpenStudy (anonymous):

Gracias :)

OpenStudy (anonymous):

welcomas

OpenStudy (anonymous):

if it didnt have = 0 would it be Ex. of a diff problem, x = 16 -4 +/- 2 sqrt3 = x + 4

OpenStudy (anonymous):

? i think you lost me

OpenStudy (anonymous):

Nevermind lol

OpenStudy (anonymous):

k

OpenStudy (anonymous):

can yu help me with this... how you would use the quadratic formula to solve x2 + 5x = –2. Why is the quadratic formula the best method to use?

OpenStudy (anonymous):

it is the best method to use because if you complete the square you are going to get annoying fractions

OpenStudy (anonymous):

solve \[x^2+5x+2=0\] via \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=5,c=2\]

OpenStudy (anonymous):

if you see \[x^2+6x-1=0\] complete the square because half of 6 is 3

OpenStudy (anonymous):

but if you see \[x^2+7x-2=0\] might as well use the formula because half of 7 is 7/2 and it will be kind of messy

OpenStudy (anonymous):

i am off. good luck!

OpenStudy (anonymous):

:)